# A contact topological glossary for non-equilibrium thermodynamics

### v2

@@ 1 Introduction
One natural question that can be addressed by the tools of contact topology is the existence of an ultrafast processes that would connect the equilibria of two thermodynamic systems. As we discuss in Theorem 5.2 below, if two equilibrium thermodynamic Legendrians are related by a slow temperature non-decreasing global process, there is an ultrafast process which starts at an equilibrium state of the first system and ends at an equilibrium state of the second system.
@@
It should be mentioned that when the process is not quasi-static (that is, if we cannot assume that at each time moment the system is in the equilibrium), the temperature of the system is not well defined from the physical point of view. We always assume that the system exchanges heat with a thermostat (or reservoir) which has a well defined temperature, and for a system not in an equilibrium, we will consider the temperature of the reservoir at the given time. In an equilibrium, the temperature of the system is equal to the temperature of the thermostat (see [27], formula (44)). In physical terms, we tacitly assume that on the “boundary of the system” the temperature equals to that of the thermostat, while “inside” the temperature is defined only in an equilibrium. In this paper, we mostly consider the temperature as an external parameter and ignore its relaxation to the equilibrium. Elimination of the entropy and the temperature from the thermodynamic phase space is called reduction, see Sections 2.3 and 3.2.
@@
Organization of the paper. Section 2 discusses the construction of the equilibrium Legendrians starting with the microscopic models and passing to the thermodynamic phase space. We also introduce the reduction process that allows us to formally freeze certain intensive and extensive thermodynamic variables. In Section 3 we explain the relation between the first and second laws of thermodynamics and the positivity of the thermodynamic paths in phase space. The definitions of the slow, fast and ultrafast processes are discussed in Section 4. Section 5 uses some non-trivial results in symplectitc topology to obtain certain consequences of the existence of slow global processes connecting the equilibrium Legendrian submanifolds for two given thermodynamic systems. Finally, in Section 6, we discuss the ultrafast processes in the examples of the ideal gas and Stirling engine, and the Curie-Weiss magnet. Interestingly, for the Stirling engine, these ultrafast processes seem to be also observed experimentally [26].
Acknowledgment. We thank Shin-itiro Goto for very useful comments on the manuscript. Preliminary results of this paper were presented by L.P. at the Symplectic Geometry Seminar in IAS, Princeton. The recording of this talk, called “Contact topology meets thermodynamics”, is available on YouTube. ME was partially supported by the Israel Science Foundation grant 2033/23, LP was partially supported by the Israel Science Foundation grant 1102/20, and LR was partially supported by NSF grant DMS-2205497 and by ONR grant N00014-22-1-2174.
@@ 2.1 Microscopic and macroscopic thermodynamic states
[…]
@@
$H(q,m)=V_{int}(m)+V_{ext}(q,m),$ (3)
where $V_{int}(m)$ is the microscopic internal energy, and $V_{ext}(q,m)$ is related to an external influence on the system, as indicated by its dependence on the exterior parameter $q\in{\mathbb{R}}^{n}$. A special case of interest is when $V(q,m)$ is linear in $q$:
$V_{ext}(q,m)=(q\cdot\overline{V}(m)).$ (4)
Here, ()($\cdot$)( ⋅ ) denotes the standard inner product on ${\mathbb{R}}^{n}$, and $\overline{V}$ is a map from $M$ to ${\mathbb{R}}^{n}$.
@@
[…]
@@
[…]
@@ 2.2 A geometric interpretation of the equilibria and the extended thermodynamic phase space
These basic notions can be represented geometrically as follows. Given a macroscopic state $\rho\in{\mathcal{P}}$, we consider the point $(-G(T,q,\rho),S(\rho),T,p(q,\rho),q)$ as an element of ${\mathbb{R}}^{2n+3}$ with the coordinates $z=-G$, $S,T$, and $(p_{j},q_{j})$, $j=1,\ldots,n$. This space is called the extended thermodynamic phase space and is denoted by $\widehat{\mathcal{T}}$. We will refer to $T$ and $q=(q_{1},\ldots,q_{n})$ as the intensive variables, and to $S$ and $p=(p_{1},\ldots,p_{n})$ as the extensive variables.
@@
called the Gibbs 1-form. It is a contact form: the maximal dimension of an integral submanifold $\Lambda$ of the distribution $\text{ker}(\lambda)$ equals $n+1$. Recall that $\Lambda$ is called integral if it is tangent to the distribution, that is, $\lambda$ vanishes on $T\Lambda$. The integral submanifolds of the maximal dimension $n+1$ are called Legendrian.
@@
$\Lambda_{f}=\Big{\{}(z,S,T,p,q,z)\in{{\mathbb{J}}^{1}{\mathbb{R}}^{n+1}}\Big{|}~{}z=f(T,q),~{}S=\frac{\partial{f}}{\partial{T}}(T,q),~{}p_{j}=\frac{\partial f}{\partial q_{j}}(T,q),j=1,\ldots,n\Big{\}}\;.$ (11)
@@
$\Psi:{\mathbb{R}}^{n+1}\times E\to{\mathbb{R}}\;,$
where $E$ is a space of auxiliary variables $\xi$, the set $\Lambda_{\Psi}$ of all $(z,S,T,p,q)\in{\mathbb{J}}^{1}{\mathbb{R}}^{n+1}$ such that there exists $\xi\in E$ so that
$z=\Psi(T,q,\xi),~{}\frac{\partial\Psi}{\partial\xi}(T,q,\xi)=0,~{}S=\frac{\partial\Psi}{\partial T}(T,q,\xi),~{}p=\frac{\partial\Psi}{\partial q}(T,q,\xi),$ (12)
is a (possibly singular) Legendrian submanifold for a generic $\Psi$. The function $\Psi$ is called a generating function of $\Lambda_{\Psi}$, and the variables $\xi$ are called the ghost variables.
A key observation is that (5)-(7) exactly mean that a Gibbs measure $\rho_{G}(T,q,\cdot)$ of a thermodynamic system corresponds to a point on the Legendrian submanifold $\Lambda$ of the form (12) with the elements $\rho\in{\mathcal{P}}$ being the ghost variables $\xi$. The generating function of $\Lambda$ is the (minus) free energy $-\Phi$ (we learned this observation from [23]). In other words, the equilibria of a thermodynamic system form a Legendrian submanifold $\widehat{\Lambda}$ in $\widehat{\mathcal{T}}$, as we vary the external parameters $T$ and $q$, that themselves are coordinates on $\widehat{\mathcal{T}}$.
@@ 2.3 The reduced thermodynamic phase space
This is an $n+2+k$ dimensional sub-space. The reduced thermodynamic phase space $\mathcal{T}$ is obtained by the projection of $h$ along the directions corresponding to the variables $S$, $p_{i},i\in I$ and $q_{e},e\in E$, and is $12k+$ 1 dimensional.
@@ 3.1 The thermodynamic paths in the extended thermodynamic phase space
To derive the aforementioned constraint, let $(T(t),q(t)),\rho(t))$ be a smooth path and $\dot{\gamma}(t)$ and $\dot{\rho}(t)$ be, respectively, the tangent vectors to the corresponding thermodynamic path $\gamma(t)$ in (15) and to $\rho(t)\in{\mathcal{P}}$. By the first law of thermodynamics, the infinitesimal increment of the macroscopic internal energy defined in (9) can be written as the difference
@@ 3.2 The thermodynamic paths in the reduced thermodynamic phase space
Let us illustrate inequality (20) in some situations. Generally, if the temperature $T$ is the only reduced variable, then $A\geq 0$ whenever the temperature is non-decreasing. This is because we always have $S(t)\geq 0$ by the definition (1) of the entropy. Similarly, in the ideal gas example considered in Section 6.1 below, the volume $V$ and the negative pressure $-P$ are pair-wise conjugate extensive and intensive variables, respectively, and the volume is, of course, always non-negative. Thus, if we reduce the pressure only, the assumption above reads that the pressure is non-increasing. On the other hand, in the Curie-Weiss model discussed in Example 3.3 and Section 6.2 below, the magnetization $M$ (dual to the external magnetic field $H$ that is an intensive variable) is an extensive variable that is not necessarily positive. In that context, when the external magnetic field is reduced, assumption (20) does not directly translate into its monotonicity in time.
@@
$0\leq\hat{\lambda}\Big{(}\frac{d\widehat{\gamma}}{dt}(t)\Big{)}$ $=\dot{z}(t)-S(t)\dot{T}(t)-\sum_{i=j}^{n}p_{j}(t)\dot{q}_{j}(t)$ (23)
@@ 4 Classification of thermodynamic processes
As we have already stated in the introduction, the main theme of this note is modeling of the non-equilibrium thermodynamic processes. We distinguish between three types of irreversible processes in the reduced (at least with respect to temperature) thermodynamic phase space. We adopt Assumption 3.2.
I. Slow (quasi-static) global processes: They are given by non-negative families (or paths) of Legendrian (equilibrium) submanifolds. Mathematically, this means that we have an initial Legendrian submanifold $\Lambda\subset\mathcal{T}$ and a map
@@
such that for each $t$, the image of $\Gamma(\cdot,t)$ is a Legendrian submanifold – such a map $\Gamma$ is called a (parameterized) Legendrian isotopy, see Section 5 below222 As we already stated in the introduction, by a global process we mean a process defined at all macroscopic equilibrium states of the system. In the geometric language, this means that the map $\Gamma$ describing this process is defined on the direct product of the whole equilibrium Legendrian submanifold $\Lambda$ – and not only of its part – with the time interval. All Legendrian submanifolds that we consider are assumed to be closed subsets of the thermodynamic phase space.. Furthermore, we assume that each individual path $t\mapsto\Gamma(x,t)$ (for a fixed $x$) is non-negative, that is, a thermodynamic path satisfying (19). (Such a Legendrian isotopy is called non-negative, see Section 5 below). Physically, this means that the change is so slow that, at each moment in time, the system is in equilibrium. An example is provided by the Curie-Weiss magnet whose temperature is slowly increasing, yielding a path of submanifolds $\Lambda(T,0)$ from Example 3.3.
A thermodynamic path $t\mapsto\Gamma(x,t)$ appearing in a slow (quasi-static) global process $\Gamma$ will be called a slow thermodynamic path.
@@
We claim that $\gamma(t)$ necessarily non-negative. Indeed, recalling (6) and since $q$ is kept fixed, we get
@@
An example is given by the thermodynamic path corresponding to the transition from the metastable to the stable equilibrium of the Curie-Weiss magnet, see [12] for a discussion of metastability and further references.
III. Ultrafast (instantaneous) processes: They correspond to Reeb chords, i.e., to pieces of trajectories of the Reeb vector field $\partial/\partial z$ connecting the initial and the terminal equilibrium Legendrian submanifolds; see [12] for a detailed discussion as well as Section 6 below. Both variables $(p,q)$ remain constant along such a process, while $z$ is increasing.
@@
As explained in [12], if the endpoint of the Reeb chord corresponds to an unstable or metastable equilibrium of the terminal system, there will be further relaxation to the stable equilibrium. We do not address the issue of the stability of equilibria in the context of Reeb chords, and it would be interesting to understand if contact topology can be helpful here. For simplicity, we stick to the terminology in III.

### v3

@@ 1 Introduction
One natural question that can be addressed by the tools of contact topology is the existence of an ultrafast processes that would connect the equilibria of two thermodynamic systems. As we discuss in Theorem 5.2 below, if two equilibrium thermodynamic Legendrians are related by a slow temperature non-decreasing global process, there is a free-energy non-increasing ultrafast process which starts at an equilibrium state of the first system and ends at an equilibrium state of the second system.
@@
It should be mentioned that when the process is not quasi-static (that is, if we cannot assume that at each moment of time the system is in the equilibrium), the temperature of the system is not well defined from the physical point of view. We always assume that the system exchanges heat with a thermostat (or reservoir) which has a well defined temperature, and for a system not in an equilibrium, we will consider the temperature of the reservoir at the given time. In an equilibrium, the temperature of the system is equal to the temperature of the thermostat (see [35], formula (44)). In physical terms, we tacitly assume that on the “boundary of the system” the temperature equals to that of the thermostat, while “inside” the temperature is defined only in an equilibrium. In this paper, we mostly consider the temperature as an external parameter and ignore its relaxation to the equilibrium. Elimination of the entropy and the temperature from the thermodynamic phase space is called reduction, see Sections 2.3 and 3.2.
@@
Organization of the paper. Section 2 discusses the construction of the equilibrium Legendrians starting with the microscopic models and passing to the thermodynamic phase space. We also introduce the reduction process that allows us to formally freeze certain intensive and extensive thermodynamic variables. In Section 3 we explain the relation between the first and second laws of thermodynamics and the positivity of the thermodynamic paths in phase space. The definitions of the slow, fast and ultrafast processes are discussed in Section 4. Section 5 uses some non-trivial results in symplectic topology to obtain certain consequences of the existence of slow global processes connecting the equilibrium Legendrian submanifolds for two given thermodynamic systems. Finally, in Section 6, we discuss the ultrafast processes in the examples of the ideal gas and Stirling engine, and the Curie-Weiss magnet. Interestingly, for the Stirling engine, these ultrafast processes seem to be also observed experimentally [34].
Acknowledgment. We thank Bernard Derrida, Leonid Levitov, Shin-itiro Goto, and Nikita Nekrasov for very useful comments. Preliminary results of this paper were presented by LP at the Symplectic Geometry Seminar in IAS, Princeton. The recording of this talk, called “Contact topology meets thermodynamics”, is available on YouTube. The paper was written during LP’s sabbatical at the University of Chicago. LP thanks the Department of Mathematics for an exceptional research atmosphere. ME was partially supported by the Israel Science Foundation grant 2033/23, LP was partially supported by the Israel Science Foundation grant 1102/20, and LR was partially supported by NSF grant DMS-2205497 and by ONR grant N00014-22-1-2174.
@@ 2.1 Microscopic and macroscopic thermodynamic states
A microscopic state of a thermodynamic system is a point on a manifold $M$, equipped with a measure $\mu$. We let $\mathcal{P}$ be the collection of smooth positive densities $\rho>0$ on $M$ so that $\int_{M}\rho d\mu=1$, and refer to an element $\rho\in\mathcal{P}$ as a macroscopic state. Its entropy is
@@
$H(q,m)=V_{\rm int}(m)+V_{\rm ext}(q,m),$ (3)
where $V_{\rm int}(m)$ is the microscopic internal energy, and $V_{\rm ext}(q,m)$ is related to an external influence on the system, as indicated by its dependence on the exterior parameter $q\in{\mathbb{R}}^{n}$. A special case of interest is when $V(q,m)$ is linear in $q$:
$V_{\rm ext}(q,m)=(q\cdot\overline{V}(m)).$ (4)
Here, $(\cdot)$ denotes the standard inner product on ${\mathbb{R}}^{n}$, and $\overline{V}$ is a map from $M$ to ${\mathbb{R}}^{n}$.
@@
${\rho_{G}:=\hbox{argmin}_{\rho\in{\mathcal{P}}}G(T,q,\rho)}=\frac{e^{-\beta H(q,m)}}{\mathcal{Z}(q)},~~\mathcal{Z}(q)=\int_{M}e^{-\beta H(q,m)}d\mu(m).$ (5)
@@
$p_{j}{(q,\rho)}=-\frac{\partial G}{\partial q_{j}}{(T,q,\rho)}=-\int_{M}\frac{\partial H}{\partial q_{i}}{(q,m)}\rho{(m)}d\mu({m})\;,\;j=1,\dots,n\;.$ (7)
@@ 2.2 A geometric interpretation of the equilibria and the extended thermodynamic phase space
These basic notions can be represented geometrically as follows. Given a macroscopic state $\rho\in{\mathcal{P}}$, we consider the point $(-G(T,q,\rho),S(\rho),T,p(q,\rho),q)$ as an element of ${\mathbb{R}}^{2n+3}$ with the coordinates z=Gz=-G, $S,T$, and $(p_{j},q_{j})$, $j=1,\ldots,n$. This space is called the extended thermodynamic phase space and is denoted by $\widehat{\mathcal{T}}$. We will refer to $T$ and $q=(q_{1},\ldots,q_{n})$ as the intensive variables, and to $S$ and $p=(p_{1},\ldots,p_{n})$ as the extensive variables.
@@
[…]
@@
$\widehat{\Lambda}_{f}=\Big\{(z,S,T,p,q,z)\in{{\mathbb{J}}^{1}{\mathbb{R}}^{n+1}}\Big|~z=f(T,q),~S=\frac{\partial{f}}{\partial{T}}(T,q),~p_{j}=\frac{\partial f}{\partial q_{j}}(T,q),j=1,\ldots,n\Big\}\;.$ (11)
@@
$\Psi:{\mathbb{R}}^{n+1}\times{\mathcal{A}}\to{\mathbb{R}}\;,$
where $\mathcal{A}$ is a linear space of auxiliary variables $\xi$, the set $\Lambda_{\Psi}$ of all $(z,S,T,p,q)\in{\mathbb{J}}^{1}{\mathbb{R}}^{n+1}$ such that there exists $\xi\in\mathcal{A}$ so that
$\widehat{\Lambda}_{\Psi}=$ $\Big\{\hbox{$(z,S,T,p,q):\exists\xi\in\mathcal{A}$ s.t }z=\Psi(T,q,\xi),~\frac{\partial\Psi}{\partial\xi}(T,q,\xi)=0,$~ (12)
$S=\frac{\partial\Psi}{\partial T}(T,q,\xi),~p=\frac{\partial\Psi}{\partial q}(T,q,\xi)\Big\},$
is a (possibly singular) Legendrian submanifold for a generic $\Psi$. The function $\Psi$ is called a generating function of $\widehat{\Lambda}_{\Psi}$, and the variables $\xi$ are called the ghost variables.
A key observation is that (5)-(7) exactly mean that a Gibbs measure $\rho_{G}(T,q,\cdot)$ of a thermodynamic system corresponds to a point on the Legendrian submanifold $\Lambda$ of the form (12) with the elements $\rho\in{\mathcal{P}}$ being the ghost variables $\xi$. The generating function of $\Lambda$ is the (minus) free energy $\Phi$ (we learned this observation from [27]). In other words, the equilibria of a thermodynamic system form a Legendrian submanifold $\widehat{\Lambda}$ in $\widehat{\mathcal{T}}$, as we vary the external parameters $T$ and $q$, that themselves are coordinates on $\widehat{\mathcal{T}}$.
@@ 2.3 The reduced thermodynamic phase space
This is an $n+2+k$ dimensional sub-space. The reduced thermodynamic phase space $\mathcal{T}$ is obtained by the projection of $h$ along the directions corresponding to the variables $S$, $p_{i},i\in I$ and $q_{e},e\in E$, and is $2k+1$ dimensional.
Remark 2.1.
Note that our notions of extended and reduced thermodynamic phase spaces and of reduction differ from the notions with the same names appearing in some other papers on geometric approaches to thermodynamics - see e.g. [36].
@@ 3.1 The thermodynamic paths in the extended thermodynamic phase space
Let us emphasize that the definition does not depend on the choice of the specific contact form $\widehat{\lambda}$ but only on the contact distribution and its co-orientation. To derive the aforementioned constraint, let $(T(t),q(t)),\rho(t))$ be a smooth path and $\dot{\gamma}(t)$ and $\dot{\rho}(t)$ be, respectively, the tangent vectors to the corresponding thermodynamic path $\gamma(t)$ in (15) and to $\rho(t)\in{\mathcal{P}}$. By the first law of thermodynamics, the infinitesimal increment of the macroscopic internal energy defined in (9) can be written as the difference
@@ 3.2 The thermodynamic paths in the reduced thermodynamic phase space
Let us illustrate inequality (20) in some situations. Generally, if the temperature $T$ is the only reduced variable, then $A\geq 0$ whenever the temperature is non-decreasing. This is because we always have $S(t)\geq 0$ by the definition (1) of the entropy. Similarly, in the ideal gas example considered in Section 6.1 below, the volume $V$ and the negative pressure P-P are pair-wise conjugate extensive and intensive variables, respectively, and the volume is, of course, always non-negative. Thus, if we reduce the pressure only, the assumption above reads that the pressure is non-increasing. On the other hand, in the Curie-Weiss model discussed in Example 3.3 and Section 6.2 below, the magnetization $M$ (dual to the external magnetic field $H$ that is an intensive variable) is an extensive variable that is not necessarily positive. In that context, when the external magnetic field is reduced, assumption (20) does not directly translate into its monotonicity in time.
@@
$0\leq\hat{\lambda}\Big(\frac{d\widehat{\gamma}}{dt}(t)\Big)$ $=\dot{z}(t)-S(t)\dot{T}(t)-\sum_{j=1}^{n}p_{j}(t)\dot{q}_{j}(t)$ (23)
@@ 4 Classification of thermodynamic processes
For our discussion on non-equilibrium thermodynamics, we have to deal with systems depending on an external parameter from a subset of a Euclidean space.
For instance, we always reduce the entropy and the temperature, considering the latter to be a parameter. As far as the other external parameters $\alpha$ are concerned, we assume that the Hamiltonian depends on $\alpha$ in the following way:
$(q,m,\alpha)\mapsto V^{\prime}_{\rm int}(m,\alpha)+q\cdot\overline{V}(m)\;,$
i.e., that only the internal energy $V^{\prime}_{\rm int}$ depends on $\alpha$.
Here is an important example. It is convenient to introduce an auxiliary external parameter to describe a jump of an extensive variable. Take a Hamiltonian $H$ of the form (3)-(4) and write a new Hamiltonian, depending on $\alpha$, as
$H(q+\alpha,m)=V^{\prime}_{\rm int}(m,\alpha)+q\cdot\overline{V}(m)=(V_{\rm int}(m)+\alpha\cdot\overline{V}(m))+q\cdot\overline{V}(m),$
where $V_{\rm int}(m)$ is the internal energy term in the original $H$ and $V^{\prime}_{\rm int}(m,\alpha):=V_{\rm int}(m)+\alpha\cdot\overline{V}(m)$ is the new internal energy depending on $\alpha$. With this notation, the jump $q\to q+c$ can be described as the jump of $\alpha$ from 0 to $c$. We refer to Remark 6.2 for further discussion.
Sometimes, we consider genuine external parameters, such as the exchange parameter of the Curie-Weiss magnet, see Section 6.3 below.
We introduce several types of non-equilibrium processes. They differ by their duration - slow (quasi-static), fast, and ultrafast (instantaneous), and by their “spread” in the thermodynamic phase space - local and global.
I. Slow (quasi-static) processes. They are caused by a slow change of some parameters of the system. The change is so slow that, at each moment in time, the system is in equilibrium. We adopt Assumption 3.2, meaning that the decrement of the free energy due to the change of the reduced intensive variables is non-negative. In particular, when the temperature and the entropy are the only reduced variables, we assume that the temperature is non-decreasing. Thus, the slow processes are given by non-negative families (or paths) of Legendrian (equilibrium) submanifolds. Mathematically, this means that we have an initial Legendrian submanifold $\Lambda\subset\mathcal{T}$ and a map
@@
such that for each $t$, the image of $\Gamma(\cdot,t)$ is a Legendrian submanifold – such a map $\Gamma$ is called a (parameterized) Legendrian isotopy, see Section 5 below. In the geometric language, this means that the map $\Gamma$ describing this process is defined on the direct product of the whole equilibrium Legendrian submanifold $\Lambda$ – and not only of its part – with the time interval. We assume that each individual path $t\mapsto\Gamma(x,t)$ (for a fixed $x$) is non-negative, that is, a thermodynamic path satisfying (24). Such a Legendrian isotopy is called non-negative, see Section 5 below. A thermodynamic path $t\mapsto\Gamma(x,t)$ appearing in a slow (quasi-static) process $\Gamma$ will be called a slow thermodynamic path.
We distinguish between two kinds of slow processes. We call a slow process global if the isotopy $\Gamma$ is a proper embedding (i.e., preimage of every compact subset is compact).  [^2] As we already stated in the introduction, by a global process we mean a process defined at all macroscopic equilibrium states of the system. An example is provided by the Curie-Weiss magnet whose temperature is slowly increasing, yielding a path of submanifolds $\Lambda(T,0)$ from Example 3.3. Otherwise we call a slow process local.
Next, we pass to a discussion of fast and ultrafast (instantaneous) processes. They are triggered by an abrupt change of the external parameters of the system, say $\alpha_{0}\to\alpha_{1}$. Assume that the initial system was in the equlibrium with the Gibbs distribution $\rho_{G,\alpha_{0}}$. The corresponding point in the thermodynamic phase space is $(z^{(0)},p^{(0)},q)$ with $z^{(0)}=-G(q,\rho_{G,\alpha_{0}},\alpha_{0})$ and
$p^{(0)}=-\int_{M}\overline{V}(m)\rho_{G,\alpha_{0}}d\mu\;.$ (27)
What happens to the system next can be described as follows.
Two stage scenario:
Stage 1: Jump. The macroscopic parameters of the system after the jump are $(z^{(1)},p^{(1)},q)$ with $z^{(1)}=-G(q,\rho_{G,\alpha_{0}},\alpha_{1})$ and
$p^{(1)}=-\int_{M}\overline{V}(m)\rho_{G,\alpha_{0}}d\mu=p^{(0)},$ (28)
cf. (27). It follows that only the free energy changes. Note that the system remains in the macroscopic state $\rho_{G,\alpha_{0}}$, while the microscopic internal energy instantaneously changes from $V_{\rm int}(m,\alpha_{0})$ to $V_{\rm int}(m,\alpha_{1})$.
Stage 2: The Fokker-Planck evolution. It belongs to the class of fast processes which we are going to discuss it right now.
@@
We claim that $\gamma(t)$ is necessarily non-negative. Indeed, recalling (6) and since $q$ is kept fixed, we get
@@
One can slightly generalize the notion of the fast process by allowing not only the temperature, but also the external parameter $\alpha(t)$ to move with time.
Let us also emphasize that here we do not assume that at time moment 0 the system is in the equilibrium. For instance, at Stage 2 above the evolution starts after the jump, and hence in general out of equilibrium. Another example is given by the thermodynamic path corresponding to the transition from the metastable to the stable equilibrium of the Curie-Weiss magnet, see [13] for a discussion of metastability and further references.
III. Ultrafast (instantaneous) processes: In the thermodynamic limit of certain mean field systems the two stage scenario above may end already at Stage 1. This happens when the limiting and the terminal Gibbs distributions coincide, so the fast process forming Stage 2 is void: the whole process consists of a jump only. Assume furthermore that these Gibbs distributions are given by a delta function on the space of microstates:
$\rho_{G,\alpha_{0}}=\rho_{G,\alpha_{1}}=\delta(m-m_{*})\;.$ (31)
The latter assumption can be verified for the Curie-Weiss magnet, and for the ideal gas. It holds true for a more general class of models, essentially due to a version of the large deviation principle. This will be discussed in a forthcoming paper.
By (27) and (28) these assumptions would yield $p:=p^{(0)}=p^{(1)}=\overline{V}(m^{*})$. Note now that the points $(p,q,z^{(0)})$ and $(p,q,z^{(1)})$ are endpoints of the Reeb chord (i.e., a segment parallel to the $z$-axes) connecting the equilibrium Legendrians.
@@
The free energy can change either way in the course of an ultrafast process. If it decreases, i.e., $z_{1}>z_{0}$, there exists a local slow process consisting of germs of Legendrians “sliding” along the chord. This happens, for instance, with the isochoric piece of the Stirling engine. It can be realized by a slow local isochoric process in the quasitatic limit (i.e., when the duration of the cycle tends to infinity), cf. [34].
Remark 4.2.
As explained in [13], if the endpoint of the Reeb chord corresponds to an unstable or metastable equilibrium of the terminal system, there will be further relaxation to the stable equilibrium. We do not address the issue of the stability of equilibria in the context of Reeb chords, and it would be interesting to understand if contact topology can be helpful here.
@@
Let us compare our description of thermodynamic processes with some earlier literature. Some authors propose to model non-equilibrium processes by contact Hamiltonian flows, see for instance [5, 17, 12, 18, 19, 20]. This approach proved successful in several interesting cases, illustrating in particular a thesis from Prigogine’s 1977 Nobel lecture [32]—that thermodynamic potentials act as Lyapunov functions near equilibrium. At the same time it has limitations in the presence of phase transitions [12]. In particular, one can prove non-existence (even locally!) of time-independent contact Hamiltonian that models some natural relaxation processes [20]. Interestingly enough, positive Legendrian isotopies, proposed as models of slow thermodynamic processes in the present paper can be realized non-uniquely by (in general) non-autonomous ambient contact Hamiltonian flows. This follows from the isotopy extension theorem in contact topology [16]. Highlighting the relevance of positive Legendrian isotopies, and in particular the link with the partial order on the space of Legendrians, is one of the main findings of the present paper. For two different contact viewpoints at the Fokker-Planck evolution (cf. the second stage of the scenario above) we refer to [13] and [19]. Let us mention also that the GENERIC formalism [21, 22] provides a simultaneous description of reversible and irreversible thermodynamic processes, including the Fokker-Planck evolution.
The interpretation of the ultrafast processes as Reeb chords between the initial and the terminal Lagrangians appeared first in [13] and is further developed in the present paper. One can analyze it through the lens of the correspondence between the quantum mechanics and thermodynamics (we thank N.Nekrasov for this suggestion). Recall that this correspondence is given by the passage to imaginary time $i\hbar/T$ which is proportional to the inverse temperature, see e.g. [14, Section 10-2] and [33, formula(2.285)]. Assume for simplicity that the Lagrangian projections $L_{0}$ and $L_{1}$ of the initial and the terminal Legendrians $\Lambda_{0}$ and $\Lambda_{1}$, respectively, are graphical of the form
$L_{j}=\Big\{p=\frac{\partial f_{j}}{\partial q}\Big\}\;,\;j=0,1\;,q\in{\mathbb{R}}^{n},$
and that they intersect transversely in a finite number of points. On the quantum side $L_{j}$ defines the space of Lagrangian (or WKB) states of the form $a_{j}(q)e^{i\frac{f_{j}(q)}{\hbar}}\;,$ with $\int|a_{j}|^{2}dq=1$, see e.g. [2]. The complex-valued functions $a_{j}$ are assumed to be smooth and compactly supported. We will denote such a state $\xi_{j}(a_{j},\hbar)$. The transition probability between $\xi_{0}(a_{0},\hbar)$ and $\xi_{1}(a_{1},\hbar)$ is given by
$p(a_{0},a_{1},\hbar):=|\langle\xi_{0}(a_{0},\hbar),\xi_{1}(a_{1},\hbar)\rangle|^{2}\;.$
Here $\langle\cdot,\cdot\rangle$ is the $L_{2}$-scalar product with respect to the Euclidean volume. The stationary phase formula (see e.g. [23, Section 15.6.2]) yields
$p(a_{0},a_{1},\hbar)=\hbar^{n/2}\int a_{0}\overline{a_{1}}\exp{\left(\frac{i}{\hbar}(f_{0}-f_{1})\right)}d\mu+O(\hbar^{1+n/2})\;,$ (32)
where $\mu$ is a complex-valued measure. Crucially, $\mu$ supported in a finite set $\mathcal{Q}$ of the $q$-coordinates of the intersection points from $L_{0}\cap L_{1}$. Note that (tautologically!) this set is in a one to one correspondence with the set of chords connecting $\Lambda_{0}$ and $\Lambda_{1}$. In fact, the points of $\mathcal{Q}$ are exactly the special values of intensive variables corresponding to ultrafast processes. It would be interesting to explore this analogy further.